# Deep Research Comparator: A Platform For Fine-grained Human Annotations of Deep Research Agents

### v1

@@ 5.2 Evaluation on Static Benchmark
To validate the performance of our baseline agent scaffold and compare human evaluation results with those from static benchmarks, we evaluated agents on DeepResearchGym Coelho et al. (2025), a benchmark that uses LLM-as-a-judge to assess long reports generated by deep research agents.
@@
Our evaluation employs two key metrics targeting different aspects of writing quality and analytical depth. We provide the LLM judge with detailed evaluation criteria and prompt it to assign scores based on it. Specifically, Clarity assesses the logical coherence and linguistic fluency of the report, while Insight captures the analytical nuance and depth of reasoning demonstrated.
5.3 Results
The annotation experiment ultimately yields 176 queries across a diverse range of domains—including science, technology, media, society, economics, health, and culture—and are complex enough to require extensive web searches to answer (See Appendix D for an example). The win rate matrix between each pair of agents is calculated based on comparison votes on these queries, as shown in Figure 5. The Bradley-Terry (BT) scores and the final ranking derived from them are presented in the first two columns of Table 2.
In addition to the side-by-side comparison votes, we collect 1,281 instances of upvote/downvote feedback on intermediate steps, and 593 instances with highlighted text spans within the reports. The final step upvote rate is shown in the third column of Table 2, and detailed statistics for fine-grained feedback are presented in Table 3.
Table 2: Evaluation results of deep research agents: human evaluation on Deep Research Comparator(Ranking, BT Score, Step Upvote Rate) and LLM-based evaluation on DeepResearchGym (Clarity, Insight).
Agent Ranking BT Score Step Upvote Rate (%) Clarity Insight
GPT Researcher 1 1135.28 74.66 89.8 90.4
Perplexity DeepResearch 2 1087.41 90.72 89.5 89.3
Simple Deepresearch (Gemini 2.5 Flash) 3 1000.00 90.33 89.4 79.4
Table 3: Upvote/downvote counts for fine-grained annotation. PD: Perplexity DeepResearch, GR: GPT Researcher, SD: Simple Deepresearch (Gemini 2.5 Flash).
Intermediate Step Text Span
Agent Upvote Downvote Upvote Downvote
GR 255 33 445 151
PR 161 17 352 36
SD 101 27 271 29
Figure 5: Win Probability of Agent A over Agent B.
The results of static benchmark evaluation for agents involved in the human annotation are presented in the last two columns of Table 2, and results for additional agents are in Table 3. Results for human agreement are presented in Table 4.
5.4 Discussion
Step-Level Feedback Analysis
Our system displays agents’ intermediate steps, with their content determined solely by agent designers. As a result, the content and granularity of steps vary across agents. Notably, although GPT Researcher ranked highest in outcome-based evaluation, it received the lowest step—upvote rate, likely due to its multi-agent parallel framework, which produces loosely ordered and often abstract or minimal step messages. In contrast, Simple Deepresearch (Gemini 2.5 Flash) present detailed steps that include both thought and actions (Appendix 2. Intermediate Steps), contributing to a higher upvote rate. This highlights the importance for agent designers to consider what step content to expose, tailoring the level of detail and structure to their annotation goals.
LLM and Human Evaluation Agreement
For report quality assessment, human rankings closely align with the LLM-based evaluations on Clarity and Insight. However, human agreement was relatively low. This may be due to the small number of queries and the fact that all three agents perform at comparably high levels (see Table 6), making fine distinctions difficult. This suggests two directions for annotation. First, as LLMs grow more capable, advancing agent performance in specialized domains may require expert annotations to provide high-quality guidance. Second, the subtlety of human preferences highlights the value of broader community annotations when the goal is to align agent behavior with general user expectations.
6 Conclusion
We present Deep Research Comparator, an open-source platform that offers a holistic framework for deep research agent hosting, web front-end demonstration, side-by-side evaluation, fine-grained annotation, and also ranking calculation. We also introduce Simple Deepresearch, a baseline scaffold that allows the easy integration of different LLMs to evaluate them as deep research agents.
@@ 6 Conclusion
Using our platform, we collected votes for 176 user queries and computed rankings based on them, and also gathered fine-grained annotation for 1,281 intermediate steps and 593 text spans in reports. Our platform can benefit a wide range of downstream applications, such as benchmarking, agent behavior analysis, and agent training via techniques like process supervision and RLHF. We believe Deep Research Comparator will serve as a useful resource for the deep research agent community.
Acknowledgements
We would like to extend our thanks Amazon AGI for providing us with the required infrastructure support for hosting Deep Research Comparator.
@@ Appendix A Simple Deepresearch Scaffold Prompt
Appendix B Unified JSON format
@@
Appendix C Fixed Question Set
  1. 1.
    How have historical trends in Spain’s real estate market influenced the development and investment potential of golf properties?
  2. 2.
    How does the inclusion of personal life details in journalist biographies affect public perception of their professional credibility and integrity?
  3. 3.
    How does product lifecycle management affect sustainability practices within the remote car starter industry?
  4. 4.
    What role does digital personalization play in enhancing user engagement with their digital devices, particularly among anime fans?
@@ 1. Text Spans
Text Span 1 “The inclusion of personal life details in journalist biographies represents a complex intersection of transparency efforts, branding strategies, and ethical considerations in contemporary journalism.” Vote: Upvote
@@
Text Span 2 “Research indicates that while personal disclosures can enhance audience familiarity with reporters, they yield minimal impact on perceptions of professional credibility and integrity.” Vote: Upvote
2. Intermediate Steps
Agent: Simple Deepresearch (Gemini Flash 2.5)
@@ 2. Intermediate Steps
Intermediate Step (Part 1) Step 1 Thought: **Planning a Deep Dive into Journalists’ Credibility** Okay, so I’m tackling this interesting question: How do the personal lives of journalists impact how the public perceives their professionalism? I’m already formulating my plan. First, I need to get organized. I’m thinking of structuring this research article in a way that builds a solid argument. My initial approach is to start with a strong **introduction**, setting the stage and clearly defining the scope of my investigation. Then, I’ll delve into a thorough **literature review**, laying out the existing studies and relevant theoretical frameworks that inform this complex relationship. Next, since this isn’t an original study, it will be a deep dive into an **analysis of existing methodology** and existing research methodologies. I need to focus on what research has already been completed, and what are its shortcomings. ACTION : Plan: 1. **Introduction:** * Define "journalist biographies" in this context (e.g., "About Me" pages, social media profiles, author bios in articles). * State the core question: How does the inclusion of personal life details affect public perception of journalist credibility and integrity? * Briefly outline the importance of the topic (e.g., trust in media, role of journalists in society). * Provide a roadmap for the article.
@@
Intermediate Step (Part 2) 2. **Theoretical Framework/Literature Review:** * **Credibility:** Define and discuss established models of source credibility (e.g., expertise, trustworthiness, goodwill). * **Integrity:** Define journalistic integrity (e.g., objectivity, fairness, accuracy, ethical conduct). * **Self-Disclosure Theory:** Explore how self-disclosure can impact perceptions and relationship building. * **Parasocial Interaction Theory:** Discuss how audiences form relationships with media figures and how personal details might foster or hinder this. * **Impression Management Theory:** How journalists might use personal details to manage their public image. * Review existing studies (if any) that touch upon personal revelations by professionals (not just journalists) and public perception. Vote: Upvote
In our preliminary data annotation experiment, the interface only had two buttons for vote: “Agent A is better” and “Agent B is better”. But in the final platform we release, all four buttons are available (“Agent A is better”, “Agent B is better”, “Tie”, and “Both are bad”).
Table 4: Human agreement results on fixed queries, measured by consensus rate. All submissions on our platform are anonymous, as the system does not require user login. Note: PD = Perplexity DeepResearch, SD = Simple Deepresearch (Gemini 2.5 Flash), GR = GPT Researcher. Green indicates the majority-voted agent, while red indicates the minority.
@@
Query GPT Researcher vs Perplexity DeepResearch Simple DeepResearch (Gemini 2.5 Flash) vs GPT Researcher Simple DeepResearch (Gemini 2.5 Flash) vs Perplexity DeepResearch
Q1 100% GR / 0% PD 60% GR / 40% SD 50% PD / 50% SD
Q2 100% GR / 0% PD 50% GR / 50% SD 80% PD / 20% SD
Q3 80% GR / 20% PD 50% GR / 50% SD 67% PD / 33% SD
Q4 75% GR / 25% PD 50% GR / 50% SD 80% PD / 20% SD
@@
Table 5: Comparison of Deep Research Comparator with related benchmarks/platforms along key dimensions.
@@
Search Arena  Miroyan et al. (2025)
Deep Research Bench   FutureSearch et al. (2025)
Deep Research Gym  Coelho et al. (2025)
BrowseComp Wei et al. (2025)
GAIA   Mialon et al. (2023)
GPQA   Rein et al. (2023)
Humanitiy’s Last Exam  HLETeam (2025)
Table 6: Scores of deep research agents on DeepResearchGym using (i) commercial search API and (ii) DeepResearchGym’s search API. Simple Deepresearch (Qwen3-8B) and Simple Deepresearch (Gemini 2.5 flash) are our baseline agent scaffold integrated with different LLMs. For GPT Researcher with commercial search API, we ran the evaluation using the newly released deep research mode, and other results are extracted from the original paper Coelho et al. (2025).

### v2

@@ 5.2. Evaluation on Static Benchmark
To compare human evaluation for overall report quality with those from static benchmarks, we evaluated agents on DeepResearchGym (coelho2025deepresearchgymfreetransparentreproducible), a benchmark that uses LLM-as-a-judge to assess reports generated by deep research agents.
@@
Our evaluation employs two metrics targeting different aspects of the overall report quality. We provide the LLM judge with detailed evaluation criteria and prompt it to assign scores based on it. Specifically, Clarity assesses the logical coherence and linguistic fluency of the report, while Insight captures the analytical nuance and depth of reasoning demonstrated.
5.3. Results
The annotation experiment yielded 176 queries across a diverse range of domains—including science, technology, society, economics, health, and culture. The resulting BT scores and the report-preference ranking are presented in Table 2. In addition to preference votes, we collected 1,281 upvote/downvote feedback instances on intermediate steps and 593 on highlighted text spans within the reports. The corresponding upvote rates are shown in Table 2.
We observe a strong alignment between our human evaluation results with static benchmark evaluation. As shown in the first four columns, the scores from the static benchmark perfectly matches the human-derived report-preference ranking. However, there’s a divergence between this overall quality based ranking and the fine-grained metric upvote rates, as shown in the last four columns.
5.4. Discussion
The strong alignment between our human annotation and static benchmark on overall report quality validates the platform’s reliability. More importantly, the divergence between this overall quality and the fine-grained metrics highlights that Deep Research Comparator provides a more nuanced perspective for agent development. The fine-grained feedback offers actionable data beyond a single outcome score. For example, text-span feedback can guide fine-grained optimization of report content, while step-wise feedback provides a valuable resource for training a process reward model or data synthesis for offline reinforcement learning.
6. Conclusion
@@
We present an open-source platform that offers a framework for holistic annotation for deep research agents. We also develop a baseline agent scaffold that allows the integration of different LLMs to evaluate them as deep research agents. Using our platform, we collected votes and fine-grained annotation for 176 user queries. We believe Deep Research Comparator will serve as a useful resource for community by benefiting a wide range of downstream applications, such as benchmarking, agent behavior analysis, and agent training via techniques like process supervision and RLHF.
7. Ethics Statement
All annotators were members of our research group who provided informed consent before contributing anonymous annotations, with the right to withdraw at any time. No personally identifiable information was collected or retained during the study.
@@ Appendix A Simple Deepresearch Scaffold Prompt
Appendix B History Context Update Mechanism of Simple Deepresearch
The history context at step $k$, $ctx_{k}$, is part of the agent’s input that contains the information of the preceding steps (from 1 to k1k-1), providing the agent with the necessary information for its current step. If $a_{k}$ is an answer action, then the generation process ends. Otherwise, the history context for step $k+1$ is updated as follows:
$ctx_{k+1}=\begin{cases}ctx_{k}+t_{k}+a_{k}+obs_{k},&a_{k}=\texttt{Search}\\ ctx_{k}+t_{k}+a_{k},&a_{k}=\texttt{Plan, Script}\\ ctx_{k},&a_{k}=\texttt{Summary}\end{cases}$
Appendix C Unified JSON format
@@ D.1. Text Spans
“The inclusion of personal life details in journalist biographies represents a complex intersection of transparency efforts, branding strategies, and ethical considerations in contemporary journalism.” Vote: Upvote
Text Span 2
@@
“Research indicates that while personal disclosures can enhance audience familiarity with reporters, they yield minimal impact on perceptions of professional credibility and integrity.” Vote: Upvote
D.2. Intermediate Steps
Agent: Simple Deepresearch
Intermediate Step (Part 1)
@@ D.2. Intermediate Steps
Step 1 Thought: **Planning a Deep Dive into Journalists’ Credibility** Okay, so I’m tackling this interesting question: How do the personal lives of journalists impact how the public perceives their professionalism? I’m already formulating my plan. First, I need to get organized. I’m thinking of structuring this research article in a way that builds a solid argument. My initial approach is to start with a strong **introduction**, setting the stage and clearly defining the scope of my investigation. Then, I’ll delve into a thorough **literature review**, laying out the existing studies and relevant theoretical frameworks that inform this complex relationship. Next, since this isn’t an original study, it will be a deep dive into an **analysis of existing methodology** and existing research methodologies. I need to focus on what research has already been completed, and what are its shortcomings. ACTION : Plan: 1. **Introduction:** * Define ”journalist biographies” in this context (e.g., ”About Me” pages, social media profiles, author bios in articles). * State the core question: How does the inclusion of personal life details affect public perception of journalist credibility and integrity? * Briefly outline the importance of the topic (e.g., trust in media, role of journalists in society). * Provide a roadmap for the article.
Intermediate Step (Part 2)
@@
2. **Theoretical Framework/Literature Review:** * **Credibility:** Define and discuss established models of source credibility (e.g., expertise, trustworthiness, goodwill). * **Integrity:** Define journalistic integrity (e.g., objectivity, fairness, accuracy, ethical conduct). * **Self-Disclosure Theory:** Explore how self-disclosure can impact perceptions and relationship building. * **Parasocial Interaction Theory:** Discuss how audiences form relationships with media figures and how personal details might foster or hinder this. * **Impression Management Theory:** How journalists might use personal details to manage their public image. * Review existing studies (if any) that touch upon personal revelations by professionals (not just journalists) and public perception. Vote: Upvote
Appendix E Detailed results for Human Annotation
Figure 3. Win Probability of Agent A over Agent B.
@@ Appendix E Detailed results for Human Annotation
Table 3. Upvote/downvote counts for fine-grained annotation. PD: Perplexity DeepResearch, GR: GPT Researcher, SD: Simple Deepresearch (Gemini 2.5 Flash).
Intermediate Step Text Span
Agent Upvote Downvote Upvote Downvote
GR 255 33 445 151
PR 161 17 352 36
SD 101 27 271 29
@@
Table 4. Comparison of Deep Research Comparator with related benchmarks/platforms along key dimensions.
@@
Search Arena  (miroyan2025searcharenaanalyzingsearchaugmented)
Deep Research Bench   (futuresearch2025deepresearchbenchevaluating)
Deep Research Gym  (coelho2025deepresearchgymfreetransparentreproducible)
BrowseComp (wei2025browsecompsimplechallengingbenchmark)
GAIA   (mialon2023gaiabenchmarkgeneralai)
GPQA   (rein2023gpqagraduatelevelgoogleproofqa)
Humanitiy’s Last Exam  (phan2025humanitysexam)
Table 5. Scores of deep research agents on DeepResearchGym using (i) commercial search API and (ii) DeepResearchGym’s search API. Simple Deepresearch (Qwen3-8B) and Simple Deepresearch (Gemini 2.5 flash) are our baseline agent scaffold integrated with different LLMs. For GPT Researcher with commercial search API, we ran the evaluation using the newly released deep research mode, and other results are extracted from the original paper (coelho2025deepresearchgymfreetransparentreproducible).
@@
Figure 4. Overview of Simple Deepresearch, our Baseline agent scaffold.